OpenStudy (anonymous):

find the valuue of k so that the graph of 21y-kx=4 and the line containing the points (5,-8) and (2,4) are parralell

OpenStudy (anonymous):

Get K by itself

OpenStudy (anonymous):

k=-(4-21y)/x

OpenStudy (anonymous):

Do you have options or is it a type in?

OpenStudy (anonymous):

typed

OpenStudy (anonymous):

In equation: 21y-kx=4, we can find the slope m1 by framing it in the form of y= mx+c ,where m is the slope 21y-kx=4 21y= 4+kx 21y= kx+4 y= kx/21+4/21 here m1= k/21 ...(1) Now slope m2 of line containing the points (5,-8) ,(2,4) m2= (y2- y1)/(x2- x1) m2= 4-(-8)/2-5 m2= 4+8/(-3) m2= -12/3 m2= -4 ...(2) Since the graph and the line are parallel, m1= m2 k/21= -4 k= -4*21 k= -84

OpenStudy (anonymous):

Sorry, you were supposed to get y by itself, not k

OpenStudy (anonymous):

so the answer is -84

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

You're welcome :)